5.6 nF, C,-8.3 nF, C,-2.1 nF) to a 9.0 volt battery. a) Determine the equivalent capacitance. b) What is the voltage across ctors are connected in parallel (C, each capacitor? c) What is the charge stored in each capacitor?
Answer #1

C1 = 5.6 nF

C2 = 8.3 nF

C3 = 2.1 nF

(a) Equivalent capacitance = C = C1 + C2 + C3 = 5.6 + 8.3 + 2.1 = 16 nF

(b) Voltage across each capacitor = V = 9.0 V

(c) Charge across C1 = Q1 = C1*V = 5.6*9 = 50.4 nC

Charge across C2 = Q2 = C1*V = 8.3*9 = 74.7 nC

Charge across C3 = Q3 = C3*V = 5.6*9 = 18.9 nC